OpenStudy (anonymous):

In studying the gene pool of a population, you find that 15% of the population is homozygous recessive (tt) for a trait. How would you use this information to solve for the frequency of the dominant T allele?

OpenStudy (anonymous):

Well if you take this question literally, you can't exactly say what's the frequency. The poppulation could have 15 % (tt) with 85 % (Tt) or 15 % (tt) with 85% (TT). Hence you must say that the chance of one individu having a dominant allele is (0,85 * 1/4) + (0,85 * 0,5) = 0,425 + 0,2125 = 0,6375 % so its around like 6 out of 10 people bring together 6 allele's with an dominant T...

OpenStudy (anonymous):

From first principles, the frequency of t (recessive) is at least 15 percent, but we need to know something about heterozygotes to be more precise. If the population is in "Hardy-Weinberg equilibrium", then we assume that the 0.15 frequency is the square of t. This is because the probability of two t alleles being in the same zygote is t^2. So that would give t = 0.39, T = 0.61.